# Data-Model Co-Evolution: Growing Test Sets to Refine LLM Behavior

### v1

@@ 7.1.3. Scaffolding Human Judgment with Just-in-Time AI Assistance
One of our design philosophy was to use AI to augment humans. The study revealed that this approach was effective. Participants found AI-generated suggestions, such as rationales for a “Bad” classification, to be particularly helpful for articulating their own reasoning, a task that can be cognitively demanding. “When I rated the AI’s response as ‘Bad’, it was helpful that it suggested reasons why. The reason for choosing ‘Bad’ can be ambiguous to put into words, but seeing the options suggested by the AI is helpful.” By offloading ancillary tasks like verbalizing a rationale, the system allowed users to focus their cognitive resources on the core challenge: making nuanced judgments.
@@ 7.2. Limitations and Design Opportunities
7.2.1. Beyond Binary: The Need for Handling Ambiguity
The system’s reliance on a binary “Good/Bad” classification can be a limitation. While simple, this binary choice forced participants into difficult decisions for borderline cases, which are often crucial for refining a model’s understanding of nuance. A participant described this struggle: “There are clear criteria for inappropriate posts, but the boundary isn’t sharp, which made evaluation feel difficult. And since there were only two choices, ‘Good’ and ‘Bad,’ for the AI’s judgment, I was conflicted on how to judge ambiguous, in-between examples.”
This is a well-known challenge in interactive machine learning (Dudley and Kristensson, 2018). As suggested by prior work (Kulesza et al., 2014; Chang et al., 2017), systems might provide users with ways to express uncertainty, such as adding a “Maybe” category. Such inputs could trigger additional components, such as seeking further clarification from the user, or explicitly modeling this uncertainty (Gordon et al., 2022).
@@ 7.2.2. The Challenge of Diverse Example Generation for Creative Exploration
This highlights a persistent challenge in generative AI: ensuring the novelty and breadth of generated content (Whitney and Norman, 2024; Reif et al., 2023). Although our system included mechanisms to encourage variety, more sophisticated strategies are needed. Future work could explore techniques for further promoting semantic diversity in generated examples, allowing users to guide the generation process toward specific types of edge cases to broaden the scope of exploration.
@@ 7.2.4. Navigating the Double-Edged Sword of Automated Judgments
The use of an LLM-as-judge (Zheng et al., 2023) presents a compelling but complex trade-off. On one hand, it can significantly reduce the burden of manual labeling and even help users articulate their thoughts. On the other hand, its fallibility can mislead users or foster over-reliance. The same participant who praised the AI’s assistance also voiced a critical concern: “After improvement, the Good/Bad predictions were often correct, but there were occasional periods where judgment became clouded. I’m concerned that some users might give up on thinking for themselves and just automate it.”
@@
7.2.5. Scaling Co-evolution: For Teams and Complex Models
Our study focused on a foundational scenario of a single application developer refining an LLM’s behavior via its prompt instruction. This setup reflects an increasingly common situation, especially in one-person startups. However, real-world applications often scale in two key dimensions: the team and the model itself. Future research could address how the co-evolution paradigm adapts to this complexity. On the human side, shifting from a solo developer to a collaborative team of product managers, policy experts, and engineers introduces significant CSCW (Computer-Supported Cooperative Work) challenges around creating shared test sets, resolving disagreements, and others (Rastogi et al., 2024; Subramonyam et al., 2021; Sivaraman et al., 2025b). Concurrently, while today’s rapid iteration happens on prompts, the “model” is a fluid concept. As techniques enabling rapid model tuning become accessible (Hu et al., 2022), the locus of iteration will shift. The core research question then includes how we can maintain the tight feedback loop of co-evolution in such environments.
@@ 8. Conclusion
In this work, we introduced the paradigm of data-model co-evolution, in which models are instantiated as prompt instructions and data as test sets, treated as co-evolving artifacts rather than separate, static resources. Through the design of an interactive workflow and its evaluation in a controlled user study, we showed that this approach enables developers to refine nuanced policies more concretely, construct richer test sets systematically, and achieve more consistent alignment between their intentions and model behavior than with a baseline prompt-editing tool. Our findings suggest that embedding iterative test-set growth into the development loop can democratize LLM refinement, moving beyond ad-hoc tinkering or episodic red-teaming toward continuous, human-in-the-loop improvement and the design of responsible, human-centric AI systems.

### v2

@@ 8.1.3. Scaffolding Human Judgment with Just-in-Time AI Assistance
One of our design philosophy was to use AI to augment humans. The study revealed that this approach was effective. Participants found AI-generated suggestions, such as rationales for an incorrect judgment, to be particularly helpful for articulating their own reasoning, a task that can be cognitively demanding. “When I rated the AI’s evaluation as ‘Incorrect’, it was helpful that it suggested reasons why. The reason for choosing ‘Incorrect’ can be ambiguous to put into words, but seeing the options suggested by the AI is helpful.”
By offloading ancillary tasks like verbalizing a rationale, the system allowed users to focus their cognitive resources on the core challenge: making nuanced judgments. This form of “just-in-time” assistance demonstrates a promising direction for human-AI collaboration.
@@ 8.2. Limitations and Design Opportunities
8.2.1. Beyond Binary Judgments: The Need for Handling Ambiguity
The system’s reliance on a binary correct/incorrect classification can be a limitation. While simple, this binary choice forced participants into difficult decisions for borderline cases, which are often the most crucial for refining a model’s understanding of nuance. A participant described this struggle: “There are clear criteria for inappropriate posts, but the boundary isn’t sharp, which made evaluation feel difficult. And since there were only two choices, ‘Correct’ and ‘Incorrect,’ for the AI’s judgment, I was conflicted on how to judge ambiguous, in-between examples.”
This is a well-known challenge in interactive machine learning (Dudley and Kristensson, 2018). As suggested by prior work (Kulesza et al., 2014; Chang et al., 2017), systems might provide users with ways to express uncertainty, such as adding a “Maybe” category. Such inputs could trigger additional components, such as seeking further clarification from the user, or explicitly modeling this uncertainty.
@@ 8.2.2. The Challenge of Diverse Example Generation for Creative Exploration
This highlights a persistent challenge in generative AI: ensuring the novelty and breadth of generated content (Whitney and Norman, 2024). Although our system included mechanisms to encourage variety, more sophisticated strategies are needed. Future work could explore techniques for further promoting semantic diversity in generated examples, allowing users to guide the generation process toward specific types of edge cases to broaden the scope of exploration.
@@ 8.2.4. Navigating the Double-Edged Sword of Automated Judgments
The use of an LLM-as-judge (Zheng et al., 2023) presents a compelling but complex trade-off. On one hand, it significantly reduces the burden of manual labeling, as discussed earlier. On the other hand, its fallibility can mislead users or foster over-reliance. The same participant who praised the AI’s assistance also voiced a critical concern: “After improvement, the Correct/Incorrect predictions were often correct, but there were occasional periods where judgment became clouded. I’m concerned that some users might give up on thinking for themselves and just automate it.”
@@
8.2.5. Extending to Open-Ended Generative Tasks.
While our user study focused on binary content moderation tasks to ensure quantifiable evaluation, the proposed workflow is conceptually applicable to open-ended generative tasks too (e.g., summarization or creative writing). If a model’s generated response is unsatisfactory, the user can still identify the failure and articulate a rationale to guide the refinement. However, we note that applying our system to generative contexts may require more sophisticated feedback mechanisms beyond binary labeling. Future work should investigate additional interaction designs to support these scenarios effectively, for instance, by integrating user-defined rubrics (Kim et al., 2024) or example demonstrations (Petridis et al., 2024).
8.2.6. Scaling to Collaborative Teams.
Our study focused on a foundational scenario of a single application developer refining an LLM’s behavior via its prompt instruction. This setup reflects an increasingly common situation, especially in one-person startups. However, real-world applications often scale to a collaborative team of product managers, policy experts, and engineers. This shifting from a solo developer to a team introduces significant CSCW (Computer-Supported Cooperative Work) challenges around creating shared test sets, resolving disagreements, and others (Rastogi et al., 2024; Subramonyam et al., 2021; Sivaraman et al., 2025b). Future research should address how the co-evolution workflow adapts to these dynamics, ensuring that the feedback loop remains tight.
8.2.7. From Prompting to Model Parameter Updates.
While our work focuses on refining prompts, which is the primary interface available to application developers today, the idea of co-evolution can be extended more broadly as a relationship between data and models. In this view, beyond updating prompt instructions, one could aim to fine-tune underlying LLMs by updating their parameters (e.g., model weights). However, unlike prompts, which can be revised rapidly, updating model parameters poses greater technical challenges. Encouragingly, recent advances in parameter-efficient fine-tuning (PEFT), including low-rank adaptation (LoRA), have begun to reduce this latency (Li and Liang, 2021; Lester et al., 2021; Hu et al., 2022). As these techniques mature, the co-evolution approach explored in our work could potentially be applied to updating model weights, increasing its practical value in settings where prompt-based control is impractical or impossible, such as under strict privacy constraints.
@@ 9. Conclusion
In this work, we introduced data-prompt co-evolution, a new workflow where prompt instructions and test data are treated as co-evolving artifacts rather than separate, static resources. Through the design of an interactive workflow and its evaluation in a controlled user study, we showed that this approach enables application developers to refine nuanced policies more concretely, construct richer test sets systematically, and achieve more consistent alignment between their intentions and model behavior than with a baseline prompt-editing tool. Our findings suggest that embedding iterative, evidence-grounded test-set growth directly into the development loop can democratize LLM refinement, moving beyond ad-hoc tinkering or episodic red-teaming toward continuous, human-in-the-loop improvement. We highlighted future opportunities for scaling co-evolution to larger test sets, collaborative settings, and model weight updates, advancing the design of responsible and human-centered AI systems.
Acknowledgements.
We thank the members of the Human-Data Interaction Lab at Yonsei University for their feedback. This work was supported in part by the Yonsei University Research Fund (2025-22-0155) and an IITP grant funded by MSIT, Korea (RS-2024-00353131).